# GUI-R1 : A Generalist R1-Style Vision-Language Action Model For GUI Agents

### v1

@@ Abstract
Existing efforts in building Graphical User Interface (GUI) agents largely rely on the training paradigm of supervised fine-tuning on Large Vision-Language Models (LVLMs). However, this approach not only demands extensive amounts of training data but also struggles to effectively understand GUI screenshots and generalize to unseen interfaces. The issue significantly limits its application in real-world scenarios, especially for high-level tasks. Inspired by Reinforcement Fine-Tuning (RFT) in large reasoning models (e.g., DeepSeek-R1), which efficiently enhances the problem-solving capabilities of large language models in real-world settings, we propose GUI-R1 , the first reinforcement learning framework designed to enhance the GUI capabilities of LVLMs in high-level real-world task scenarios, through unified action space rule modeling. By leveraging a small amount of carefully curated high-quality data across multiple platforms (including Windows, Linux, MacOS, Android, and Web) and employing policy optimization algorithms such as Group Relative Policy Optimization (GRPO) to update the model, GUI-R1 achieves superior performance using only 0.02% of the data (3K vs. 13M) compared to previous state-of-the-art methods like OS-Atlas across eight benchmarks spanning three different platforms (mobile, desktop, and web). These results demonstrate the immense potential of reinforcement learning based on unified action space rule modeling in improving the execution capabilities of LVLMs for real-world GUI agent tasks.
(a) Grounding Capability
(b) Low-level Task Capability
(c) High-level Task Capability
@@ 1 Introduction
  • We conduct a comprehensive evaluation of GUI agents, covering three distinct platforms (desktop, mobile, and web) and three levels of task granularity (low-level grounding, low-level tasks, and high-level tasks) across eight benchmarks. Experimental results demonstrate that our RUN1-R1 is leading in multiple realistic cases. This creates a strong baseline of GUI agents for future research.
@@ 2.2 Reinforcement Fine-Tuning
3 RUN1-R1 Framework
Figure 2: Overview of the GUI-R1 Framework. Given the high-level instruction , action history and visual image inputs, the policy model generates multiple responses containing reasoning steps. Then the verifiable rewards, such as action type reward, click point reward, and input text reward, are used with the policy gradient optimization algorithm to update the policy model.
@@ 3.2 Verifiable Rewards in Unified Action Space
Unified Action Space Prompt for Task Training and Inference You are RUN1-R1, a reasoning GUI Agent Assistant. In this UI screenshot $\boldsymbol{<image>}$, I want you to continue executing the command $\boldsymbol{task}$, with the action history being $\boldsymbol{history}$. Please provide the action to perform (enumerate from [complete, close/delete, press_home, click, press_back, type, select, scroll, enter]), the point where the cursor is moved to (integer) if a click is performed, and any input text required to complete the action. Output the thinking process in $<$$\boldsymbol{think}$$>$ $<$$/\boldsymbol{think}$$>$ tags, and the final answer in $<$$\boldsymbol{answer}$$>$ $<$$/\boldsymbol{answer}$$>$ tags as follows: $<$$\boldsymbol{think}$>>><<<$/\boldsymbol{think}$$>$ $<$$\boldsymbol{answer}$$>$[‘action’: enum[complete, close/delete, press_home, click, press_back, type, select, scroll, enter], ‘point’: [x, y], ‘input_text’: ‘no input text [default]’]$<$$/\boldsymbol{answer}$$>$.
Unified Action Space Prompt for Grounding Training and Inference You are RUN1-R1, a reasoning GUI Agent Assistant. In this UI screenshot $\boldsymbol{<image>}$, I want you to continue executing the command $\boldsymbol{task}$, with the action history being $\boldsymbol{history}$. Please provide the action to perform (enumerate from [click]), the point where the cursor is moved to (integer) if a click is performed, and any input text required to complete the action. Output the thinking process in $<$$\boldsymbol{think}$$>$ $<$$/\boldsymbol{think}$$>$ tags, and the final answer in $<$$\boldsymbol{answer}$$>$ $<$$/\boldsymbol{answer}$$>$ tags as follows: $<$$\boldsymbol{think}$>>><<<$/\boldsymbol{think}$$>$ $<$$\boldsymbol{answer}$$>$[‘action’: enum[click], ‘point’: [x, y], ‘input_text’: ‘no input text [default]’]$<$$/\boldsymbol{answer}$$>$.
@@ 3.3 Training Data Curation
Figure 3: Illustrations of image and action category quantity and difficulty distributions in GUI-R1 -3K.
@@ 4.3 Ablation Study
Image resolution and data quality. To investigate the impact of image resolution and data quality on GUI RFT, we conduct corresponding ablation experiments, with the results shown in Figure 4. As observed, when using the filtered RUN1-R1-3K dataset, the model requires only a few updates to achieve relatively high rewards. In contrast, training with unfiltered and low-quality data necessitates significantly more training time for the model to converge, with a noticeably lower performance ceiling. To further explore the effect of image resolution on model training, we increase the image resolution to twice its original size (from 1,048,576 pixels to 2,097,152 pixels). As shown in Figure 4, because of the high resolution of GUI task images and the small size of many UI elements, increasing the image resolution allows the model to perceive these elements more clearly, which accelerates the convergence speed of RFT and improves the performance ceiling.
Coefficients in the reward function. In order to explore the impact of the coefficients for format rewards and accuracy rewards in the reward function on the final performance, we conduct relevant ablation experiments, as shown in the Table 4. The results indicate that reducing the coefficient ratio of format rewards leads to consistent performance improvements. This is because format rewards are easier to learn during training and often converge early in the process. By amplifying the accuracy rewards, the advantages of providing correct answers are further emphasized, ultimately leading to more performance improvements.
Figure 4: Ablation Study of Image resolution and data quality
@@ 4.4 Visualization
In fig. 5, we provide additional visualization of the training process. As shown in fig. 5(a) and fig. 5(b), it can be observed that the format reward converges quickly in the early stages of training, while the accuracy reward becomes the main source of differentiated rewards in the later stages of training. Furthermore, as illustrated in fig. 5(d), the mean response length first decreases and then gradually increases, but the ”aha moment” does not occur. This may be due to the single-image input training method in a non-interactive environment, which prevents the model from autonomously tracing back the sequence of incorrect actions. Exploring multi-image high-level tasks in interactive environments could be a potential direction for inducing the emergence of the ”aha moment” in future research.
@@
Figure 5: Visualization of Training Process of GUI-R1 . To provide more details, we report the curves of GUI-R1 ’s key metrics during training, including format reward, accuracy reward, mean response length, and policy gradient (PG) loss, as they vary with the training steps.

### v3

@@ Abstract
Existing efforts in building graphical user interface (GUI) agents largely rely on the training paradigm of supervised fine-tuning (SFT) on large vision-language models (LVLMs). However, this approach not only demands extensive amounts of training data but also struggles to effectively understand GUI screenshots and generalize to unseen interfaces. The issue significantly limits its application in real-world scenarios, especially for high-level tasks. Inspired by reinforcement fine-tuning (RFT) in large reasoning models (e.g., DeepSeek-R1), which efficiently enhances the problem-solving capabilities of large language models in real-world settings, we propose GUI-R1, the first reinforcement learning framework designed to enhance the GUI capabilities of LVLMs in high-level real-world task scenarios, through unified action space rule modeling. By leveraging a small amount of carefully curated high-quality data across multiple platforms (including Windows, Linux, MacOS, Android, and Web) and employing policy optimization algorithms such as group relative policy optimization (GRPO) to update the model, GUI-R1 achieves superior performance using only 0.02% of the data (3K vs. 13M) compared to previous state-of-the-art methods like OS-Atlas across eight benchmarks spanning three different platforms (mobile, desktop, and web). These results demonstrate the immense potential of reinforcement learning based on unified action space rule modeling in improving the execution capabilities of LVLMs for real-world GUI agent tasks. The codebase is available at https://github.com/ritzz-ai/GUI-R1.git.
(a) Grounding capability.
(b) Low-level task capability.
(c) High-level task capability.
@@ 1 Introduction
  • We conduct a comprehensive evaluation of GUI agents, covering three distinct platforms (desktop, mobile, and web) and three levels of task granularity (low-level grounding, low-level tasks, and high-level tasks) across eight benchmarks. Experimental results demonstrate that our GUI-R1 is leading in multiple realistic cases. This creates a strong baseline of GUI agents for future research.
@@ 2.2 Reinforcement Fine-Tuning
3 GUI-R1 Framework
Figure 2: Overview of the GUI-R1 Framework. Given the high-level instruction, action history, and visual image inputs, the policy model generates multiple responses containing reasoning steps. Then the verifiable rewards, such as action type reward, click point reward, and input text reward, are used with the policy gradient optimization algorithm to update the policy model.
@@ 3.2 Verifiable Rewards in Unified Action Space
Unified Action Space Prompt for Task Training and Inference You are GUI-R1, a reasoning GUI Agent Assistant. In this UI screenshot $\boldsymbol{<image>}$, I want you to continue executing the command $\boldsymbol{task}$, with the action history being $\boldsymbol{history}$. Please provide the action to perform (enumerate from [complete, close/delete, press_home, click, press_back, type, select, scroll, enter]), the point where the cursor is moved to (integer) if a click is performed, and any input text required to complete the action. Output the thinking process in $<$$\boldsymbol{think}$$>$ $<$$/\boldsymbol{think}$$>$ tags, and the final answer in $<$$\boldsymbol{answer}$$>$ $<$$/\boldsymbol{answer}$$>$ tags as follows: $<$$\boldsymbol{think}$>>><<<$/\boldsymbol{think}$$>$ $<$$\boldsymbol{answer}$$>$[‘action’: enum[complete, close/delete, press_home, click, press_back, type, select, scroll, enter], ‘point’: [x, y], ‘input_text’: ‘no input text [default]’]$<$$/\boldsymbol{answer}$$>$.
Unified Action Space Prompt for Grounding Training and Inference You are GUI-R1, a reasoning GUI Agent Assistant. In this UI screenshot $\boldsymbol{<image>}$, I want you to continue executing the command $\boldsymbol{task}$, with the action history being $\boldsymbol{history}$. Please provide the action to perform (enumerate from [click]), the point where the cursor is moved to (integer) if a click is performed, and any input text required to complete the action. Output the thinking process in $<$$\boldsymbol{think}$$>$ $<$$/\boldsymbol{think}$$>$ tags, and the final answer in $<$$\boldsymbol{answer}$$>$ $<$$/\boldsymbol{answer}$$>$ tags as follows: $<$$\boldsymbol{think}$>>><<<$/\boldsymbol{think}$$>$ $<$$\boldsymbol{answer}$$>$[‘action’: enum[click], ‘point’: [x, y], ‘input_text’: ‘no input text [default]’]$<$$/\boldsymbol{answer}$$>$.
@@ 3.3 Training Data Curation
Figure 3: Illustrations of image and action category quantity and difficulty distributions in the dataset GUI-R1-3K.
@@ 4.3 Ablation Study
Image resolution and data quality. To investigate the impact of image resolution and data quality on GUI RFT, we conduct corresponding ablation experiments, with the results shown in Figure 4. As observed, when using the filtered GUI-R1-3K dataset, the model requires only a few updates to achieve relatively high rewards. In contrast, training with unfiltered and low-quality data necessitates significantly more training time for the model to converge, with a noticeably lower performance ceiling. To further explore the effect of image resolution on model training, we increase the image resolution to twice its original size (from 1,048,576 pixels to 2,097,152 pixels). As shown in Figure 4, because of the high resolution of GUI task images and the small size of many UI elements, increasing the image resolution allows the model to perceive these elements more clearly, which accelerates the convergence speed of RFT and improves the performance ceiling.
Coefficients in the reward function. To explore the impact of the coefficients for format rewards and accuracy rewards in the reward function on the final performance, we conduct relevant ablation experiments, as shown in Table 4. The results indicate that reducing the coefficient ratio of format rewards leads to consistent performance improvements. This is because format rewards are easier to learn during training and often converge early in the process. By amplifying the accuracy rewards, the advantages of providing correct answers are further emphasized, ultimately leading to more performance improvements.
Figure 4: Ablation study of image resolution and data quality.
@@ 4.4 Visualization
In Figure 5, we provide additional visualization of the training process. As shown in Figure 5(a) and Figure 5(b), it can be observed that the format reward converges quickly in the early stages of training, while the accuracy reward becomes the main source of differentiated rewards in the later stages of training. Furthermore, as illustrated in Figure 5(d), the mean response length first decreases and then gradually increases, but the “aha moment” does not occur. This may be due to the single-image input training method in a non-interactive environment, which prevents the model from autonomously tracing back the sequence of incorrect actions. Exploring multi-image high-level tasks in interactive environments could be a potential direction for inducing the emergence of the “aha moment” in future research.
@@
Figure 5: Visualization of the training process of GUI-R1 . To provide more details, we report the curves of GUI-R1 ’s key metrics during training, including format reward, accuracy reward, mean response length, and policy gradient (PG) loss, as they vary with the training steps.